A particle originally at rest at the highest point of a smooth vertical circle is slightly displaced. It will leave the circle at a vertical distance h below the highest point such that
(a)h=R (b)h=R/3 (c)h=R/2 (d)h=2R/3
(R is the radius of the circle it is moving in)

Grade:12th pass

3 Answers

Arun
5 years ago
Dear Anant
 
I am getting my answer as h = R/3
 
Please let me know is thisis correct then I will share my solution
Anant Kumar Thakur
5 years ago
yeah thats the answer please share the solution...................................................................
Khimraj
5 years ago
let angle made by particle with verticl line is \Theta
when centrifugal force (mv2/r) become greater than or equal tocomponent of mg along center (mgcos\Theta)
mv2/R >= mgcos\Theta
m(2gh)/r = mg(R-h)/R
2h = R-h
so h = R/3
Hope it clears......................